OpenStudy (anonymous):

Part 1: Solve the inequality -5x + 8 <= x - 4 and show all of your work. You can use <= to represent the less than or equal to symbol. Part 2: Use complete sentences to describe the graph of the solution from Part 1. Be sure to include the type of circle and which direction the shading will be.

OpenStudy (anonymous):

-5x + 8 <= x - 4 -5x - x + 8 <= x - x - 4 -6x + 8 <= - 4 -6x + 8 - 8 <= - 4 - 8 -6x <= - 12 -6x/-6 >= -12/-6 x >= 2 <--0------1-------2///////3//> " /// " represents shaded area. x >= 2 reads..x is greater then or equal to 2 You would put a solid circle on 2 because 2 is included in the solution. Then you would shade to the right because x is greater or equal to 2.